# Mixed Dynamics In Linear Networks: Unifying the Lazy and Active Regimes

### v1

@@ Abstract
The training dynamics of linear networks are well studied in two distinct setups: the lazy regime and balanced/active regime, depending on the initialization and width of the network. We provide a surprisingly simple unyfing formula for the evolution of the learned matrix that contains as special cases both lazy and balanced regimes but also a mixed regime in between the two. In the mixed regime, a part of the network is lazy while the other is balanced. More precisely the network is lazy along singular values that are below a certain threshold and balanced along those that are above the same threshold. At initialization, all singular values are lazy, allowing for the network to align itself with the task, so that later in time, when some of the singular value cross the threshold and become active they will converge rapidly (convergence in the balanced regime is notoriously difficult in the absence of alignment). The mixed regime is the ‘best of both worlds’: it converges from any random initialization (in contrast to balanced dynamics which require special initialization), and has a low rank bias (absent in the lazy dynamics). This allows us to prove an almost complete phase diagram of training behavior as a function of the variance at initialization and the width, for a MSE training task.
@@ 1.1 Contributions
We consider the training dynamics of shallow linear networks $A_{\theta}=W_{2}W_{1}$ and show that for large enough width $w$ (the inner dimension), and a iid $\mathcal{N}(0,\sigma^{2})$ initialization of all weights, the dynamics of $A_{\theta(t)}$ as a result of training the parameters $\theta=(W_{1},W_{2})$ with GD/GF on the loss $\mathcal{L}(\theta)=C(A_{\theta})$ for a general matrix cost $C$ is approximately given by the self-consistent dynamics
$\partial_{t}A_{\theta(t)}\approx-\sqrt{A_{\theta}A_{\theta}^{T}+\sigma^{4}w^{2}I}\nabla C(A_{\theta})-\nabla C(A_{\theta})\sqrt{A_{\theta}^{T}A_{\theta}+\sigma^{4}w^{2}I}.$ (1)
@@
$\partial_{t}A_{\theta(t)}\approx-2\sigma^{2}w\nabla C(A_{\theta})$
@@
$\partial_{t}A_{\theta(t)}=\sqrt{A_{\theta}A_{\theta}^{T}}\nabla C(A_{\theta})-\nabla C(A_{\theta})\sqrt{A_{\theta}^{T}A_{\theta}}$
@@
  • Active ($1>\gamma_{\sigma^{2}}+\gamma_{w}$) where $K=\mathrm{Rank}A^{*}$ pass the threshold and fit $A^{*}$ before the other singular values have time to fit the noise $E$, leading to the recovery of $A^{*}$.
@@ 1.2 Previous Works
Linear networks have been used as a testing ground, a stepping stone on the way to understand nonlinear networks. Linear networks and their training dynamics are many ways much simpler than nonlinear ones, but in spite of a long research history, our understanding remains limited.
The setting that is best understood is that of diagonal linear networks where the dynamics decouple along the diagonal entries leading to an incremental learning behavior and a sparsity bias saxe_2014_exact ; advani_2017_independent_diag ; Saxe_2019_independent_diag ; gidel_2019_independent_diag , some of this analysis has been extended to include effects of initialization scale woodworth_2020_diagnet_bias and SGD Pesme_2021_SGD_bias_diag_nets . While the same decoupling happens in general linear with diagonal initializations and diagonal task, it remains an extremely strong assumption.
Some work has been done to prove similar incremental learning dynamics outside the diagonal case li2020towards ; jacot-2021-DLN-Saddle ; jin_2023_incremental_learning . For shallow networks, the training dynamics with MSE can be explicitly solved Fukumizu1998_solution_shallow_DLN but remain very complex so that one needs to assume some form of alignment to guarantee convergence Braun_2022_exact_DLN . For deeper networks there exists explicit formulas in the mean-field limit where the number of neurons grows to infinity chizat_2022_mean_field_DLN .
A very powerful tool in the analysis of linear network is its training invariants, and the balancedness condition which greatly simplifies the dynamics arora_2018_DLN_convergence ; arora_2018_depth_speed_Lp . Balanced networks exhibit a momentum effect, where the training dynamics along a singular value $s_{i}$ have ‘speed’ proportional to $s_{i}$ itself (or $s_{i}$ to some power), while this momentum effect seems to be key to understand the low-rank bias of linear networks arora_2019_matrix_factorization , it also means that one needs to guarantee that the dynamics never approach zero, which is one the main hurdle towards proving convergence in balanced networks. To solve this issue, recent work has focused on initialization that slightly imbalanced xiong_2023_overparam_balanced_hurt_conv ; min_2021_imbalance_shallow_DLN ; tarmoun_2021_GF_matrix_fact ; min_2023_imbalance_deep ; xu_2023_imbalance_GD . This suggests that it is key to find the right balance between balancedness and imbalancedness to obtain both fast convergence and low-rank bias.
@@ 2 Mixed Lazy/Balanced Dynamics
For a linear net $A_{\theta}=W_{2}W_{1}$ with width $w$, initialized with i.i.d. $\mathcal{N}(0,\sigma^{2})$ weights and trained with Gradient Flow, we have with prob. $1-\delta$ for all times $t$
$\left\|C_{1}(t)-\hat{C}_{1}(t)\right\|_{op},\left\|C_{2}(t)-\hat{C}_{2}(t)\right\|_{op}$ $\leq\text{min}\left\{O(\sigma^{2}w),O\left(\sqrt{\frac{d}{w}}\left\|C_{1}\right\|_{op}\right)\right\}.$
@@
00 0 $\approx C_{2}^{3}-A_{\theta}A_{\theta}^{2}C_{2}-C_{2}A_{\theta}A_{\theta}^{T}-\sigma^{4}w^{2}C_{2}+A_{\theta}C_{1}A_{\theta}^{T}.$
@@ 2.1 Phase Diagram for MSE
We can identify a region of ‘reasonable’ pairs $(\gamma_{\sigma^{2}},\gamma_{w})$ by ruling out degenerate behavior. First, the width $w$ needs to be larger than the dimension $d$, since a network of width $w$ can only represent matrices of rank $w$ or less, this means that we need $\gamma_{w}\geq 1$. Another constraint comes from the variance of $A_{\theta}$ at initialization: the entries $A_{\theta(0),ij}$ at initialization have variance $\sigma^{4}w$. We want this variance to not blow up as $d$ grows which implies that we need $2\gamma_{\sigma^{2}}+\gamma_{w}\leq 0$.
@@ A.1 Convention and Notation
Constants. $d$ is the input dimension, $K$ is the rank of matrix $A^{*}$, and $w$ is the input and output dimension. $c$ and $C$ will usually denote constants that are independent of $d$, and depending on the context, the value of $c$ and $C$ might be different. If $x$ is a scalar that depends on $d$ and $y$ is a scalar, then $x=O(y)$ means there exists a constant $c$ independent of $d$, such that for $d$ sufficiently large we have $x\leq cy$. If $A$ is a matrix, then $A=O(y)$ means there exists a constant $c$ independent of $d$ such that for $d$ sufficiently large, $\|A\|_{op}\leq cy$. $O(y)$ can be either a matrix or a scalar, and its meaning will be always clear from context.
Matrix. We use $\|\cdot\|_{op},\|\cdot\|_{F},\|\cdot\|$ to denote operator norm of a matrix, Frobenius norm of a matrix, and the norm of vectors. For every matrix $A$, we define $\text{max}A$ as the $L^{\infty}$ norm, $\text{max}_{i,j}\absolutevalue{A_{ij}}$ |. We use $I$ to denote identity matrix, the dimension of which is determined by context. We shall assume that the signal singular values of $A^{*}$ are $s_{1}^{*},\ldots,s_{K}^{*}$. $\forall i=1,2,\ldots,K$, $s_{i}=a_{i}d$ where $a_{1}\geq a_{2}\geq\ldots\geq a_{K}$ are constants independent of $d$. By selecting proper basis in the input and output space, we assume that $A^{*}=S^{*}$, where $S^{*}$ is the diagonal matrix consisting of singular values of $A^{*}$, ordered from largest to smallest. The $p,q$-th element of a matrix $A$ is denoted $A_{pq}$. We reserve the notation $A(i,j)$ for the $i,j$-th block matrix of $A$, which we shall define below.
@@
A.2 Matrix Norms
@@
The two equations above proves the first claim.
@@ A.3 Perturbation of Singular Values and Singular Vectors
We will often use the following variant of Davis-Kahan $\sin\theta$ theorem.
@@
. Let $\Sigma,\hat{\Sigma}\in\mathbb{R}^{p\times p}$ be symmetric, with eigenvalues $\lambda_{1}\geq\ldots\lambda_{p}$ and $\hat{\lambda}_{1}\geq\ldots\hat{\lambda}_{p}$. Fix $1\leq r\leq s\leq p$, let $d=r-s+1$ and let $V=(v_{r},\ldots,v_{s})$ and $\hat{V}=(\hat{v}_{r},\ldots,\hat{v}_{s})$ have orthonormal columns satisfying $\Sigma v_{j}=\lambda_{j}v_{j}$ and $\Sigma\hat{v}_{j}=\hat{\lambda}_{j}v_{j}$. Let $\sigma_{1},\ldots,\sigma_{d}$ be the singular values of $\hat{V}^{T}V$. Let $\Theta(V,\hat{V})$ be the diagonal matrix with $\cos\Theta(V,\hat{V})_{jj}=\sigma_{j}$ and $\sin\Theta(V,\hat{V})$ be defined entry-wise. Then
@@
The term $\|\sin\Theta(V,\hat{V})\|_{F}$ describes the extend to which $V$ is aligned with $\hat{V}$, but it is a bit complicated to take derivative. In this paper we will use a different characterization of alignment.
@@
Let $\hat{\Sigma}$ be a $d\times d$ diagonal matrix, and let $s_{1},\ldots,s_{K},\ldots,s_{d}$ be its diagonal entries. Assume that $s_{1}=\ldots=s_{n_{1}}>s_{n_{1}+1}=\ldots=s_{n_{2}}\geq\ldots s_{n_{m}}=s_{K}>s_{K+1}=\ldots=s_{d}$. Then $\hat{\Sigma}$ has $m+1$ blocks in total. Assume that $\|X-\hat{\Sigma}\|_{op}=d^{\alpha}$, which is very small. Let $X=USV^{T}$, and define
@@
Let $X=USV^{T}$. Then $X^{T}X=VS^{2}V^{T}$ is a symmetric matrix. Let $V(:,k)=(v_{n_{k-1}+1},\ldots,v_{n_{k}})$ be the singular vectors of $X$ corresponding to $s_{n_{k-1}+1},\ldots,s_{n_{k}}$. There are some freedom to choose $\hat{V}$, but for simplicity we pick $\hat{V}(:,k)=(e_{n_{k-1}+1},\ldots,e_{n_{k}})$ where $e_{i}$ is the $i$-th coordinate vector. As a result, $\hat{V}(:,k)^{T}V(:,k)=V(k,k)$. Apply DK-$\sin\theta$ theorem to $V$ and $\hat{V}$, we see that
@@
In the case where $d$ increases to infinity, $m$ is a constant, $s_{1}=O(s_{K})$, $s_{K}>cs_{K+1}$, $K$ is a constant, $\|X-\hat{\Sigma}\|_{op}=o(\|\hat{\Sigma}\|_{op})$, and $\text{min}_{k}(s_{n_{k}}^{2}-s_{n_{k+1}}^{2})\geq cs_{K}^{2}$ for some constant $c$, we have $x\leq O(\frac{\|X-\hat{\Sigma}\|_{op}}{\|\hat{\Sigma}\|_{op}})$.

### v2

@@ Abstract
The training dynamics of linear networks are well studied in two distinct setups: the lazy regime and balanced/active regime, depending on the initialization and width of the network. We provide a surprisingly simple unifying formula for the evolution of the learned matrix that contains as special cases both lazy and balanced regimes but also a mixed regime in between the two. In the mixed regime, a part of the network is lazy while the other is balanced. More precisely the network is lazy along singular values that are below a certain threshold and balanced along those that are above the same threshold. At initialization, all singular values are lazy, allowing for the network to align itself with the task, so that later in time, when some of the singular value cross the threshold and become active they will converge rapidly (convergence in the balanced regime is notoriously difficult in the absence of alignment). The mixed regime is the ‘best of both worlds’: it converges from any random initialization (in contrast to balanced dynamics which require special initialization), and has a low rank bias (absent in the lazy dynamics). This allows us to prove an almost complete phase diagram of training behavior as a function of the variance at initialization and the width, for a MSE training task.
@@ 1.1 Contributions
We consider the training dynamics of shallow linear networks $A_{\theta}=W_{2}W_{1}$ and show that for large enough width $w$ (the inner dimension), and a iid $\mathcal{N}(0,\sigma^{2})$ initialization of all weights, the dynamics of $A_{\theta(t)}$ as a result of training the parameters $\theta=(W_{1},W_{2})$ with GD/GF on the loss $\mathcal{L}(\theta)=C(A_{\theta})$ for a general matrix cost $C$ with learning rate $\eta$ is approximately given by the self-consistent dynamics
$\partial_{t}A_{\theta(t)}\approx-\eta\sqrt{A_{\theta}A_{\theta}^{T}+\sigma^{4}w^{2}I}\nabla C(A_{\theta})-\eta\nabla C(A_{\theta})\sqrt{A_{\theta}^{T}A_{\theta}+\sigma^{4}w^{2}I}.$ (1)
@@
$\partial_{t}A_{\theta(t)}\approx-2\eta\sigma^{2}w\nabla C(A_{\theta})$
@@
$\partial_{t}A_{\theta(t)}=-\eta\sqrt{A_{\theta}A_{\theta}^{T}}\nabla C(A_{\theta})-\eta\nabla C(A_{\theta})\sqrt{A_{\theta}^{T}A_{\theta}}$
@@
  • Active ($1>\gamma_{\sigma^{2}}+\gamma_{w}$) where $K=\mathrm{Rank}A^{*}$ singular values pass the threshold and fit $A^{*}$ before the other singular values have time to fit the noise $E$, leading to the recovery of $A^{*}$.
@@ 1.2 Previous Works
Linear networks have been used as a testing ground, a stepping stone on the way to understand nonlinear networks. Linear networks and their training dynamics are in many ways much simpler than nonlinear ones, but in spite of a long research history, our understanding remains limited.
The setting that is best understood is that of diagonal linear networks where the dynamics decouple along the diagonal entries leading to an incremental learning behavior and a sparsity bias saxe_2014_exact ; advani_2017_independent_diag ; Saxe_2019_independent_diag ; gidel_2019_independent_diag ; Pesme_2023_S-to-S_diag_DLN , some of this analysis has been extended to include effects of initialization scale woodworth_2020_diagnet_bias and SGD Pesme_2021_SGD_bias_diag_nets . While the same decoupling happens in general linear with diagonal initializations and diagonal task, it remains an extremely strong assumption.
Some work has been done to prove similar incremental learning dynamics outside the diagonal case li2020towards ; jacot-2021-DLN-Saddle ; jin_2023_incremental_learning where the incremental aspect can be understood as the parameters going from saddle to saddle. For shallow linear networks, the training dynamics with MSE can be explicitly solved Fukumizu1998_solution_shallow_DLN but remain very complex so that one needs to assume some form of alignment to guarantee convergence Braun_2022_exact_DLN . For deeper networks there exists explicit formulas in the mean-field limit where the number of neurons grows to infinity chizat_2022_mean_field_DLN , these results can of course be applied to the special case of shallow nets, our paper goes further by giving self-consistent dynamics for the full matrix, revealing the lazy/active transition, and also extends the analysis to finite widths.
A very powerful tool in the analysis of a linear network is its training invariants, and the balancedness condition which greatly simplifies the dynamics arora_2018_DLN_convergence ; arora_2018_depth_speed_Lp . Balanced networks exhibit a momentum effect, where the training dynamics along a singular value $s_{i}$ have ‘speed’ proportional to $s_{i}$ itself (or $s_{i}$ to some power), while this momentum effect seems to be key to understand the low-rank bias of linear networks arora_2019_matrix_factorization , it also means that one needs to guarantee that the dynamics never approach zero, which is one the main hurdle towards proving convergence in balanced networks. To solve this issue, recent work has focused on initialization that slightly imbalanced xiong_2023_overparam_balanced_hurt_conv ; min_2021_imbalance_shallow_DLN ; tarmoun_2021_GF_matrix_fact ; min_2023_imbalance_deep ; xu_2023_imbalance_GD . This suggests that it is key to find the right balance between balancedness and imbalancedness to obtain both fast convergence and low-rank bias.
In a concurrent work kunin_2024_richquickexactsolutions a similar transition between lazy and active regimes is observed, and the same mixed dynamics are derived for a specific initialization. In contrast, we prove that these dynamics are approximately true with high probability for random Gaussian initializations, which is the standard initialization scheme for neural networks.
@@ 2 Mixed Lazy/Balanced Dynamics
For a linear net $A_{\theta}=W_{2}W_{1}$ with width $w$, initialized with i.i.d. $\mathcal{N}(0,\sigma^{2})$ weights and trained with Gradient Flow, we have with high probability that for all time $t$,
$\left\|C_{1}(t)-\hat{C}_{1}(t)\right\|_{op},\left\|C_{2}(t)-\hat{C}_{2}(t)\right\|_{op}$ $\leq\text{min}\left\{O(\sigma^{2}w),O\left(\sqrt{\frac{d}{w}}\left\|C_{1}(t)\right\|_{op}\right)\right\}.$
@@
00 0 $\approx C_{2}^{3}-A_{\theta}A_{\theta}^{T}C_{2}-C_{2}A_{\theta}A_{\theta}^{T}-\sigma^{4}w^{2}C_{2}+A_{\theta}C_{1}A_{\theta}^{T}.$
@@
The takeaway from theorem 1 is the following.
  1. 1.
    In the lazy regime where $\|C_{1}(t)\|_{op}+\|C_{2}(t)\|_{op}\leq O(\sigma^{2}w)$, then $\|C_{1}(t)-\hat{C}_{1}(t)\|_{op}\leq\sqrt{d/w}\|C_{1}(t)\|_{op}<<\|C_{1}(t)\|_{op}$.
  2. 2.
    In the active regime where $\|C_{1}(t)\|_{op}/\sigma^{2}w>d^{\varepsilon}>>1$, then $\|C_{1}(t)-\hat{C}_{1}(t)\|_{op}\leq O(\sigma^{2}w)<<d^{-\varepsilon}\|C_{1}(t)\|_{op}$.
It is true that the error does not vanish. However, for our purpose it suffices to show that $\|\hat{C}_{1}-C_{1}\|_{op}$ is infinitely smaller than $C_{1}$ for all times, regardless of the magnitude of $\|C_{1}(t)\|_{op}$.
@@ 2.1 Phase Diagram for MSE
We can identify a region of ‘reasonable’ pairs $(\gamma_{\sigma^{2}},\gamma_{w})$ by ruling out degenerate behavior. First, the width $w$ needs to be larger than the dimension $d$, since a network of width $w$ can only represent matrices of rank $w$ or less, this means that we need $\gamma_{w}\geq 1$. Another constraint comes from the variance of $A_{\theta}$ at initialization: the entries $A_{\theta(0),ij}$ at initialization have variance $\sigma^{4}w$. We want this variance to go to zero as $d$ grows which implies that we need $02\gamma_{\sigma^{2}}+\gamma_{w}<$ 0.
@@ A.1 Convention and Notation
Constants. $d$ is the dimension of the input and the output layer, $K$ is the rank of matrix $A^{*}$, and $w$ is the dimension of hidden layer. $c$ and $C$ will usually denote constants that are independent of $d$, and depending on the context, the value of $c$ and $C$ might be different. If $x$ is a scalar that depends on $d$ and $y$ is a scalar, then $x=O(y)$ means there exists a constant $c$ independent of $d$, such that for $d$ sufficiently large we have $x\leq cy$. If $A$ is a matrix, then $A=O(y)$ means there exists a constant $c$ independent of $d$ such that for $d$ sufficiently large, $\|A\|_{op}\leq cy$. $O(y)$ can be either a matrix or a scalar, and its meaning will be always clear from context.
Matrix. We use $\|\cdot\|_{op},\|\cdot\|_{F},\|\cdot\|$ to denote the operator norm of a matrix, the Frobenius norm of a matrix, and the $L^{2}$ norm for vectors. For every matrix $A$, we define $\text{max}A$ as the $L^{\infty}$ norm, $\text{max}_{i,j}\absolutevalue{A_{ij}}$ |. We use $I$ to denote identity matrix, the dimension of which is determined by context. We shall assume that the signal singular values of $A^{*}$ are $s_{1}^{*},\ldots,s_{K}^{*}$. $\forall i=1,2,\ldots,K$, and $s_{i}^{*}=a_{i}d$ where $a_{1}\geq a_{2}\geq\ldots\geq a_{K}$ are constants independent of $d$. By selecting proper basis in the input and output space, we assume that $A^{*}=S^{*}$, where $S^{*}$ is the diagonal matrix consisting of singular values of $A^{*}$, ordered from largest to smallest. The $p,q$-th element of a matrix $A$ is denoted $A_{pq}$. We reserve the notation $A(i,j)$ for the $i,j$-th block matrix of $A$, which we shall define below.
@@
Element-wise Product.We use $\odot$ to represent element-wise product of two matrix of the same shape.
Important Assumptions. Throughout the paper, we shall always assume that
  1. 1.
    $\gamma_{w}>1$ (i.e., $w>>d$).
  2. 2.
    $02\gamma_{\sigma^{2}}+\gamma_{w}<$ 0. (i.e., $\sigma^{4}w<<1$).
A.2 Matrix Inequalities
In the proof of main theorems, we will work extensively with inequalities of matrix norms and inequalities that involves element-wise product. The element-wise product appears naturally in the derivative of singular vectors of a matrix.
@@ A.2 Matrix Inequalities
The two equations above prove the first claim.
@@ A.3 Perturbation of Singular Values and Singular Vectors
We will often use the following variant of the Davis-Kahan $\sin\theta$ theorem.
@@
. Let $\Sigma,\hat{\Sigma}\in\mathbb{R}^{p\times p}$ be symmetric, with eigenvalues $\lambda_{1}\geq\ldots\geq\lambda_{p}$ and $\hat{\lambda}_{1}\geq\ldots\geq\hat{\lambda}_{p}$. Fix $1\leq r\leq s\leq p$, let $d=r-s+1$ and let $V=(v_{r},\ldots,v_{s})$ and $\hat{V}=(\hat{v}_{r},\ldots,\hat{v}_{s})$ have orthonormal columns satisfying $\Sigma v_{j}=\lambda_{j}v_{j}$ and $\Sigma\hat{v}_{j}=\hat{\lambda}_{j}v_{j}$. Let $\sigma_{1},\ldots,\sigma_{d}$ be the singular values of $\hat{V}^{T}V$. Let $\Theta(V,\hat{V})$ be the diagonal matrix with $\cos\Theta(V,\hat{V})_{jj}=\sigma_{j}$ and $\sin\Theta(V,\hat{V})$ be defined entry-wise. Then
@@
The implication of the theorem is that if two matrices are sufficiently close, then their singular vectors are also close to each other. In the case where $r=s$ and $\lambda_{r}$ is of multiplicity 1, the theorem reduces to saying that the sine value of the angle between $v_{r}$ and $\hat{v}_{r}$ is very small.
The term $\|\sin\Theta(V,\hat{V})\|_{F}$ is complicated to take derivative. In this paper we will use the following characterization of alignment, which is easier to take derivatives.
@@
Let $\hat{\Sigma}$ be a $d\times d$ diagonal matrix, and let $s_{1},\ldots,s_{K},\ldots,s_{d}$ be its diagonal entries. Assume that $s_{1}=\ldots=s_{n_{1}}>s_{n_{1}+1}=\ldots=s_{n_{2}}\geq\ldots s_{n_{m}}=s_{K}>s_{K+1}=\ldots=s_{d}$. Then $\hat{\Sigma}$ has $m+1$ blocks in total. Assume that $\|X-\hat{\Sigma}\|_{op}=d^{\alpha}$. Let $X=USV^{T}$, and define
@@
Let $X=USV^{T}$. Then $X^{T}X=VS^{2}V^{T}$ is a symmetric matrix. Let $V(:,k)=(v_{n_{k-1}+1},\ldots,v_{n_{k}})$ be the singular vectors of $X$ corresponding to $s_{n_{k-1}+1},\ldots,s_{n_{k}}$. There is some freedom to choose $\hat{V}$, but for simplicity we pick $\hat{V}(:,k)=(e_{n_{k-1}+1},\ldots,e_{n_{k}})$ where $e_{i}$ is the $i$-th coordinate vector. As a result, $\hat{V}(:,k)^{T}V(:,k)=V(k,k)$. Apply DK-$\sin\theta$ theorem to $V$ and $\hat{V}$, we see that
@@
In the case where $d\to\infty$ , $m$ is a constant, $\alpha<1$, $s_{1}=O(s_{K})$, $s_{K}>cs_{K+1}$, $K$ is a constant, $\|X-\hat{\Sigma}\|_{op}=o(\|\hat{\Sigma}\|_{op})$, and $\text{min}_{k}(s_{n_{k}}^{2}-s_{n_{k+1}}^{2})\geq cs_{K}^{2}$ for some constant $c$, we have $x\leq O(\frac{\|X-\hat{\Sigma}\|_{op}}{\|\hat{\Sigma}\|_{op}})$.